# Teaching LLMs According to Their Aptitude: Adaptive Reasoning for Mathematical Problem Solving

### v3

@@ Abstract
Xin Xu1, Yan Xu2, Tianhao Chen1, Yuchen Yan3, Chengwu Liu4, Zaoyu Chen5, Yufei Wang2, Yichun Yin2, Yasheng Wang2, Lifeng Shang2, Qun Liu2 1The Hong Kong University of Science and Technology, 2Huawei Noah’s Ark Lab, 3Zhejiang University, 4Peking University, 5The Hong Kong Polytechnic University xxuca@connect.ust.hk, xuyan110@huawei.com
@@ 3.3 Implicit Instruction Tuning
[…]
@@ 4.1 Problem Setting
Figure 2: Overview of our Teaching LLMs According to Their Aptitude (TATA) framework. Here, $\mathcal{D}_{\text{orig}}$ denotes the original training set, $\mathcal{D}_{\text{aug}}$ represents the augmented training set obtained through rejection sampling with CoT only, and $\mathcal{D}$ refers to the candidate set consisting of (query, CoT, TIR) triplets. $\mathcal{D}_{\text{anchor}}$ is the anchor set of size $A$. ${S_{\text{CoT}}^{k}}$ and $S_{\text{TIR}}^{k}$ are scores calculated based on the LLMs’ aptitude on the anchor set, elicited using 1-shot prompts. Finally, $\mathcal{H}$ represents the SFT data selection process. Fine-tuning on the resulting SFT data enables LLMs to spontaneously select between CoT and TIR at test time according to their aptitude.
@@ 4.3 TATA Details
Currently, two scores, ${S_{\text{CoT}}^{k}}$ and $S_{\text{TIR}}^{k}$, are associated with the $k$-th query $q_{k}$ in the candidate set $\mathcal{D}$. The next step is to determine whether to include the CoT or the TIR solutions for this specific query $q_{k}$ in $\mathcal{D}$. Specifically, the goal is to decide between $\{(x_{k},y_{k}^{j})\}_{j=1}^{M_{k}}\subseteq D_{\text{SFT}}$ or $\{(x_{k},z_{k}^{j})\}_{j=1}^{M_{k}}\subseteq D_{\text{SFT}}$. We formalize this decision process with a decision function $\mathcal{H}_{k}=(S_{\text{CoT}}^{k},S_{\text{TIR}}^{k})$, where the final decision is represented as a series of decisions =$\mathcal{H}=\{\mathcal{H}_{k}\}_{k=1}^{N}$, where $N$ is the number of queries in candidate set $\mathcal{D}$. For instance, a simple decision function $\mathcal{H}_{k}$ could involve consistently choosing CoT solutions, i.e., $\{(x_{k},y_{k}^{j})\}_{j=1}^{M_{k}}\subseteq D_{\text{SFT}}$ for all $k$. This corresponds to performing SFT exclusively on CoT data.
@@ A.2 TIR Inference Pipeline
1:problem $q$, model $\mathcal{G}$, prompt ${p}$, external tools $\mathcal{E}$, stop condition Stop($\cdot$), maximum iteration rounds $n$
2:$\tau_{0}\leftarrow\text{""}$" $\triangleright$ Trajectory Initialization
@@
4:     $r_{i}\sim\mathbb{P}_{\mathcal{G}}(\cdot|p\oplus q\oplus\tau_{i-1})$triangleright Rationale Generation
5:     if Stop($r_{i}$then $\triangleright$ Stopping Criteria
@@
8:     $a_{i}\sim\mathbb{P}_{\mathcal{G}}(\cdot|p\oplus q\oplus\tau_{i-1}\oplus r_{i})$triangleright Program Generation
9:     $o_{i}\leftarrow\mathcal{E}(a_{i})$triangleright Tool Execution
10:     $\tau_{i}\leftarrow\tau_{i-1}\oplus r_{i}\oplus a_{i}\oplus o_{i}$triangleright Trajectory Update
@@ A.3 Implicit Instruction Tuning
[…]
@@ B.1 TATA Implementation Details
[…]

### v4

@@ Abstract
Xin Xu1, Yan Xu2, Tianhao Chen1, Yuchen Yan3, Chengwu Liu4, Zaoyu Chen5, Yufei Wang2, Yichun Yin2, Yasheng Wang2, Lifeng Shang2, Qun Liu2, Lu Yin6 1The Hong Kong University of Science and Technology, 2Huawei Noah’s Ark Lab, 3Zhejiang University, 4Peking University, 5The Hong Kong Polytechnic University, 6University of Surrey xxuca@connect.ust.hk, xuyan110@huawei.com
@@ 3.3 Implicit Instruction Tuning
where $\|$ denotes concatenation. The first term only involves the test input $\mathbf{X_{\text{test}}}$, and the second term is related to few-shot exemplars, which can be interpreted as an IIT to the model parameters (Dai et al., 2022; Yang et al., 2023) (see Appendix A.3).
@@ 4.1 Problem Setting
Figure 2: Overview of our Teaching LLMs According to Their Aptitude (TATA) framework. Here, $\mathcal{D}_{\text{orig}}$ denotes the original training set, $\mathcal{D}_{\text{aug}}$ represents the augmented training set obtained through rejection sampling with CoT only, and $\mathcal{D}$ refers to the candidate set consisting of (query, CoT, TIR) triplets. $\mathcal{D}_{\text{anchor}}$ is the anchor set of size AA. $S_{\text{CoT}}^{k}$} and $S_{\text{TIR}}^{k}$ are scores calculated based on the LLMs’ aptitude on the anchor set, elicited using 1-shot prompts. Finally, $\mathcal{H}$ represents the SFT data selection process. Fine-tuning on the resulting SFT data enables LLMs to spontaneously select between CoT and TIR at test time according to their aptitude.
@@ 4.3 TATA Details
Currently, two scores, $S_{\text{CoT}}^{k}$} and $S_{\text{TIR}}^{k}$, are associated with the $k$-th query $q_{k}$ in the candidate set $\mathcal{D}$. The next step is to determine whether to include the CoT or the TIR solutions for this specific query $q_{k}$ in $\mathcal{D}$. Specifically, the goal is to decide between $\{(x_{k},y_{k}^{j})\}_{j=1}^{M_{k}}\subseteq D_{\text{SFT}}$ or $\{(x_{k},z_{k}^{j})\}_{j=1}^{M_{k}}\subseteq D_{\text{SFT}}$. We formalize this decision process with a decision function $\mathcal{H}_{k}=(S_{\text{CoT}}^{k},S_{\text{TIR}}^{k})$, where the final decision is represented as a series of decisions $\mathcal{H}=\{\mathcal{H}_{k}\}_{k=1}^{N}$, where $N$ is the number of queries in candidate set $\mathcal{D}$. For instance, a simple decision function $\mathcal{H}_{k}$ could involve consistently choosing CoT solutions, i.e., $\{(x_{k},y_{k}^{j})\}_{j=1}^{M_{k}}\subseteq D_{\text{SFT}}$ for all $k$. This corresponds to performing SFT exclusively on CoT data.
@@ A.2 TIR Inference Pipeline
1:problem $q$, model $\mathcal{G}$, prompt p{p}, external tools $\mathcal{E}$, stop condition Stop($\cdot$), maximum iteration rounds $n$
[…]
@@
4:  $r_{i}\sim\mathbb{P}_{\mathcal{G}}(\cdot|p\oplus q\oplus\tau_{i-1})$ $\triangleright$ Rationale Generation
5:  if Stop($r_{i}$) then $\triangleright$ Stopping Criteria
@@
8:  $a_{i}\sim\mathbb{P}_{\mathcal{G}}(\cdot|p\oplus q\oplus\tau_{i-1}\oplus r_{i})$ $\triangleright$ Program Generation
9:  $o_{i}\leftarrow\mathcal{E}(a_{i})$ $\triangleright$ Tool Execution
10:  $\tau_{i}\leftarrow\tau_{i-1}\oplus r_{i}\oplus a_{i}\oplus o_{i}$ $\triangleright$ Trajectory Update
@@ A.3 Implicit Instruction Tuning
In-Context Learning (ICL) can be interpreted as a form of implicit instruction tuning, where the model is effectively "fine-tuned" using the given demonstrations in an implicit manner (Dai et al., 2022; Yang et al., 2023; Irie et al., 2022; Li et al., 2023). Let $\mathbf{X_{\text{ins}}},\mathbf{X_{\text{test}}}\in\mathbb{R}^{d_{\text{in}}}$ represent the few-shot demonstration inputs and the test input, respectively. We define the attention query vector as $\textit{{Q}}=\mathbf{W}_{Q}\mathbf{X_{\text{test}}^{\top}}$, while the attention key and value vectors are given by $\textit{{K}}=\mathbf{W}_{K}[\mathbf{X}_{\text{ins}}\|\mathbf{X}_{\text{test}}]$ and $\textit{{V}}=\mathbf{W}_{V}[\mathbf{X}_{\text{ins}}\|\mathbf{X}_{\text{test}}]$, where $\|$ denotes concatenation. The projection matrices $\mathbf{W}_{K},\mathbf{W}_{V},\mathbf{W}_{Q}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ are used to compute the attention queries, keys, and values. The self-attention mechanism for a single attention head in any given layer is formulated as follows:
@@ B.1 TATA Implementation Details
2. Present the final result in using a $\boxed{\text{ANS}}$ without any units.